Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
lumaca, nardecchia, casanueva, gherardini - 9:37 Thursday 24 October 2024 (65385) Print this report
WI DAS IN tuning

Since the locking issues during the night (65381) and still ongoing, the foreseen activity of TCS tunining (WI DAS INner ring power to be rised up: 154 mW --> 300 mW) is post postponed until the interferometer is working properly again.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×